OpenStudy (anonymous):

Indefinite integrals? Not sure how different they are from regular integrals. I'm really bad at anti derivatives so if they come up can someone explain how you found the antiderivative as well? integral of 6x^2(2x^3 - 5)^5 dx

OpenStudy (anonymous):

substitute 2x^3-5 = t

OpenStudy (anonymous):

6x^2 dx = dt

OpenStudy (anonymous):

\[\int_{10}^{\infty}\frac{6x^2}{(2x^3-5)^5}dx\]?? you said it was "improper" but did not specify the limits of integration

OpenStudy (anonymous):

I don't know what you mean satellite73

OpenStudy (anonymous):

oh nvm you said "indefinite" ignore me i don't know how to read

OpenStudy (anonymous):

what @divu.mkr said, put \(u=3x^2-5, du =6x^2dx\) and you are done in one step

OpenStudy (anonymous):

Since it is indefinite and there are no limits to fix, that's all I need to finish the problem?

OpenStudy (freckles):

and he meant u=2x^3-5 :p

OpenStudy (anonymous):

write all in terms of \(u\)

OpenStudy (anonymous):

yeah that one \(u=2x^3-5\) sorry

OpenStudy (anonymous):

Ah okay. Thank you very much! Teacher gave me a packet on Friday and I've been slowly working through them for a couple days.

OpenStudy (freckles):

you can do the integration part right? not the sub part I guess? I'm just asking so I know what part is the part you find difficult

OpenStudy (anonymous):

I think so. I'm working through it right now to see what I get

OpenStudy (freckles):

post here if you want it checked

OpenStudy (anonymous):

Okay I got (2x^3 - 5)^6 / 6

OpenStudy (freckles):

+C and yah

OpenStudy (anonymous):

uggggghhh I always forget +C Alright thanks to those who replied!